6. An automobile whose speed is increasing at a rate of 0.6 m/s2

 travels along a circular road of

radius 20 m. When the instantaneous speed of the automobile is 4 m/s, find (a) the tangential

acceleration component, (b) the centripetal acceleration component


1
Expert's answer
2021-07-21T09:06:12-0400

(a)

"a_t=0.6\ \dfrac{m}{s^2}."


(b)

"a_c=\dfrac{v^2}{r}=\dfrac{(4\ \dfrac{m}{s})^2}{20\ m}=0.8\ \dfrac{m}{s^2}."